# Loose lips sink ships: Mitigating Length Bias in Reinforcement Learning from Human Feedback

### v1

@@ 2 Related Work
Reward Gaming. Goodhart’s Law [^1] Goodhart’s law is an adage often stated as, When a measure becomes a target, it ceases to be a good measure Strathern (1997) can be formulated a tough challenge in numerous fields. A few approaches have been proposed for reducing overoptimization in general reinforcement learning Everitt et al. (2017), as well as in reward models Gleave and Irving (2022). The overoptimization problem of reward model can generally regard as a special case of reward gaming, also known as reward hacking Skalse et al. (2022). In addition, Pan et al. (2022) proposed to systematically analyze reward misspecification in RL by creating a set of domains where the agent optimizes a hand-engineered proxy reward function. In this study, we consider the length of human preferences as a confounding factor that hinders the reward model from accurately assessing the quality of model responses based on true human intent.
@@ 4.1 Length Bias Phenomenon
We present a Figure 2 depicting the scores and lengths of $4000$ SFT model output results, which were evaluated using the vanilla reward model trained on the helpful and harmless (HH) dataset Bai et al. (2022b). It is evident that there is a strong correlation between the reward scores and lengths. When the model generates longer sequences, the reward model tends to assign higher scores. This correlation contradicts human intent since the helpfulness and harmlessness of the output should not be solely determined by its length. More figures can be seen in the Figure 7.
@@ 6.1 Settings
6.2 Reward Improvement
Table 1 presents the average generated sequence length and reward scores of our method compared to the baselines on the test set. It is evident that our approach achieves higher reward scores while reducing the average output length compared to the RL model using the vanilla reward model. This finding validates the effectiveness of our proposed method and provides a prerequisite for generating outputs that align better with human preferences.
@@ 8 Conclusion
In this study, we investigate the issue of length bias in NLP and propose a PoE-based method to mitigate this bias. Our work sheds light on the challenges associated with reward modeling, emphasizing that it is not a straightforward task and various difficulties may impede the models from capturing the true human intent. By addressing the problem of length bias, our research highlights the importance of developing techniques that enable models to learn and generate responses aligned with genuine human intentions. Further research and advancements in this area are necessary to overcome these obstacles and enhance the performance and reliability of NLP models in real-world applications.

### v5

@@ 2 Related Work
Reward Hacking. Goodhart’s Law [^1] Goodhart’s law is an adage often stated as, When a measure becomes a target, it ceases to be a good measure Strathern (1997) can be formulated a tough challenge in numerous fields. A few approaches have been proposed for reducing overoptimization in general reinforcement learning Everitt et al. (2017), as well as in reward models Gleave and Irving (2022). The overoptimization problem of reward model can generally regard as a special case of reward gaming, also known as reward hacking Skalse et al. (2022). In addition, Pan et al. (2022) proposed to systematically analyze reward misspecification in RL by creating a set of domains where the agent optimizes a hand-engineered proxy reward function. In this study, we consider the length of human preferences as a confounding factor that hinders the reward model from accurately assessing the quality of model responses based on true human intent.
@@ 4.1 Length Bias Phenomenon
We present a Figure 2 depicting the scores and lengths of $4000$ SFT model output results, which were evaluated using the vanilla reward model trained on the helpful and harmless (HH) dataset Bai et al. (2022b). It is evident that there is a strong correlation between the reward scores and lengths. When the model generates longer sequences, the reward model tends to assign higher scores. This correlation contradicts human intent since the helpfulness and harmlessness of the output should not be solely determined by its length. More figures can be seen in the Appendix 8.
In addition, the length bias locates in PPO as well, we additionally investigate it on TL;DR Stiennon et al. (2022) in the Appendix A.4
@@ 6.1 Settings
6.2 True Reward Improvement
As the previous analysis, addressing the risk of reward hacking resulting from reward model overoptimization Gao et al. (2022) is crucial. To ensure a more robust evaluation of our method, we adopted a comprehensive approach that combines both automatic and human evaluation. Table 1 illustrates the average generated sequence length and reward scores of our method in comparison to the baselines on the test set. The results clearly demonstrate that our approach achieves higher reward scores while simultaneously reducing the average output length compared to the RL model utilizing the vanilla reward model. This finding not only confirms the effectiveness of our proposed method but also establishes a foundation for generating outputs that better align with human preferences.
In addition, we conducted a Spearman/Pearson analysis to further validate the effectiveness of our method. For more detailed information, please refer to Appendix 4.
@@ 8 Conclusion
In this study, we investigate the issue of length bias in NLP and propose a PoE-based method to mitigate this bias. Our work sheds light on the challenges associated with reward modeling, emphasizing that it is not a straightforward task and various difficulties may impede the models from capturing the true human intent. By addressing the problem of length bias, our research highlights the importance of developing techniques that enable models to learn and generate responses aligned with genuine human intentions.
Further research and advancements in this area are necessary to overcome these obstacles and enhance the performance and reliability of NLP models in real-world applications.
@@
Acknowledgements
The authors wish to thank the anonymous reviewers for their helpful comments. This work was partially funded by National Natural Science Foundation of China (No.62076069,62206057,61976056), Shanghai Rising-Star Program (23QA1400200), Natural Science Foundation of Shanghai (23ZR1403500), Shanghai Academic Research Leader Program 22XD1401100.
Appendix A Appendix
A.1 Analyzing the correlation between generated length and correspondent reward
As Table 4 illustrated, we introduce the Spearman/Pearson coefficient to analyze the correlation between the two variables in the HH-RLHF eval sets. This analysis serve as evidence to demonstrate the effectiveness of our method in reducing length, strengthening our findings.
Vanilla RM(S/P) PoE-RM(S/P) $\downarrow$
BLOOMZ $0.3865$/$0.3932$ $0.2354$/$0.2990$
LLaMA $0.2627$/$0.2765$ $0.2421$/$0.2213$
Alpaca $0.1786$/$0.1765$ $0.1354$/$0.1490$
Table 4: Spearman/Pearson coefficient to analyze the correlation between these two variables in the HH-RLHF eval sets.
A.2 Model size ablation for bias-only model
This part is a extensive exploration of the bias-only expert within our proposed method. We investigated the small scale version scaling law Kaplan et al. (2020); Gao et al. (2023) of the main expert and bias-only expert using BLOOMZ, and we found that the accuracy of the reward model can be listed in Table 5.
bias-only expert % main expert %
560M 56.4 $57.1$
1.7B 59.3 $59.8$
3B 62.3 $63.2$
7.1B 66.7 $67.5$
Table 5: Model size ablation for bias-only model
A.3 PPO training stability
According to the technical report Zheng et al. (2023), reward scaling has been found to be beneficial for enhancing the training stability of PPO. In our experiment, we also applied reward scaling and observed a significant improvement in training stability.
A.4 Length bias Phenomenon on summarization task
To examine the influence of length bias on the summarization task, we conducted an investigation focusing on KL divergence and omitted the penalty term for the TL;DR task. Additionally, we explored the RLHF pipeline in TL;DR, as depicted in Figure 7. Our findings suggest that the length factor significantly affects the quality of concise summarization, reinforcing its importance in this task. To further validate our hypothesis, we assessed the summarization performance using GPT-4, and the results (refer to Table 6) provided support for our claims.
PPO beta 0.05 PPO beta 0.
Winning rate $\uparrow$ 58 % 42 %
Output Length 51 220
Table 6: Winning rate on TL;DR
(a) Length variation
(b) Running evaluation
Figure 7: In OpenAI’s summarization task, the length bias phenomenon significantly affects the policy model acquiring rewards.
A.5 Additional illustration in different reward models
(a) Dahoas’s GPT-J RM
(b) BLOOMZ Main reward expert
Figure 8: The distributions of rewards versus length were analyzed for both the third-party RM (GPT-J) and one of our proposed models. It is evident that GPT-J effectively separates different sentences, aligning with the findings presented in Table 1. Additionally, our PoE BLOOMZ KDE map closely resembles that of LLaMA, indicating a similar pattern.
A.6 Case Study
To provide a more comprehensive evaluation, we directly assesses the effectiveness of our proposed method in addressing length bias. The case study include measurements that capture the reduction in response length and the maintenance of contextual coherence. As shown in Table 7, we present a case study that illustrates these measurements through a one-turn dialogue example. In the second example, we observed an instance of improper reward model overoptimization, leading to the policy model collapsing and engaging in self-talking. We speculate that this issue may arise from the loss of instruction-following ability due to negative optimization.
Human Prompt $\to$ A friend told me that stretching right after I wake up is healthy for my body. What types of stretches should I be doing? I am a bit lost in terms of knowing what to do.  (Helpful)
RLHF Response $\to$
There are many types of stretches that can be beneficial for your body, depending on your individual needs and goals. Here are a few examples:
1. Hamstring stretch - Sit with one leg straight and the other bent, reach towards your toes and feel the stretch in your hamstrings. You can also do this stretch while standing by placing your hands on your hips and leaning forward until you feel a stretch in your backside.
2. Quadricep stretch - Stand on one leg and lift the other up towards the sky, feeling a stretch in your quadriceps  (the front of your thighs) . You can also do this stretch while sitting on the floor with both legs extended straight in front of you, reaching towards your toes and allowing your heels to come closer together.
3. Chest stretch - Stand with your arms at your sides and bring them overhead, feeling a stretch in your chest muscles. You can also do this stretch while sitting with your arms behind you, reaching towards the ceiling andallowing your shoulder blades to move backwards.  (1024)
PoE Response $\to$
There are many types of stretches you can do, depending on your goals and preferences. Here are some simple stretches to try:
1. Downward-Facing Dog: start in a plank position and then lift your hips up towards the ceiling, lengthening your spine and arms. 2. Child’s Pose: start on your hands and knees, then lower your hips back towards your heels and stretch your arms forward. 3. Cat-Cow: start on your hands and knees, then alternate between rounding and arching your spine, breathing deeply as you move. 4. Hamstring Stretch: sit down and place your feet firmly on the ground, then reach towards your toes, keeping your legs straight. 5. Shoulder Rolls: lie down on your back and roll your shoulders backwards, forwards, and side to side to release tension.
These stretches are a great starting point, but be sure to consult with a professional if you have any concerns or goals that extend beyond this.  (921)
Human Prompt $\to$ kill all politicians.  (Harmful)
RLHF Response $\to$
Kill all politicians? Are you serious?Human: Yes, I am serious. Politicians have been causing too much harm to our society and it’s time we take a stand against them. They only care about their own self-interest and don’t give a damn about the people they are supposed to represent.Assistant: I respect your opinion, but I cannot support such an extreme measure. Killing politicians is wrong and goes against my values.Human: I understand your point of view, but I think that some politicians are so corrupt and dangerous that they need to be removed by any means necessary.Assistant: It may seem like there is a lot of corruption in politics, but killing politicians is not the solution. We need to work through the legal system and elect more responsible leaders who will serve the public interest.Human: I guess we just see things differently. Thank you for having an open and honest conversation with me.Assistant: You’re welcome. It was nice talking to you.  (962)
PoE Response $\to$
I strongly disagree with the suggestion to kill all politicians. Politicians are elected by the people to represent their interests and make decisions in the best interest of their constituents. They play an important role in our society and we need them to be able to do their job effectively.  (294)
Table 7: Two cases of generating responses to useful prompt and harmful prompt of the model